# Detecting and Evaluating Medical Hallucinations in Large Vision Language Models

### v1

@@ 6.3 Evaluations of the MediHallDetector’s Performance
Comparison of different hallucination detection models: We compared our model’s performance with two of the most powerful LLMs: GPT-3.5 and GPT-4 in the context of medical hallucination hierarchy. As depicted in the radar chart in Figure 2(d), the chart illustrates the alignment of detection models with human preferences across six different hallucination levels. GPT-3.5 predominantly classified responses as Critical and Attribute hallucinations, and only correctly identified approximately 1.14% of these instances, indicating a poor understanding of the nuanced definitions of different hallucination levels and an inability to reasonably assess the correctness of $A$ in more complex hallucination classification scenarios. While GPT-4 performed better at detecting the correctness of $A$ in complex prompts, they still struggled with the hierarchical categorization of hallucinations. This was particularly evident in distinguishing prompt-induced hallucinations.
Figure 2(c) compares MedHallDetector with GPT-3.5 and GPT-4 across two dimensions: evaluation time and multi-round evaluation consistency. The results show that GPT models tend to produce varying conclusions in each evaluation, whereas MedHallDetector maintains a high consistency rate of 93.88% across three evaluation rounds. Additionally, MedHallDetector achieves the shortest average inference time per evaluation.
@@ 7 Conclusion
In this paper, we address the challenges of hallucination detection and evaluation in the application of LVLMs in healthcare. We propose a novel benchmark, evaluation metrics, and detection model specifically designed for the medical domain. In addition to establishing baselines for current mainstream models on the benchmark, we demonstrate the effectiveness of our metrics and model in hallucination evaluation and detection through extensive experimental analysis. We hope this work can significantly improve the reliability of LVLMs in medical applications.
@@ A.1 Details of Med-HallMark
Conventional Medical Questions: In this paper, conventional medical questions refer to the original questions we constructed based on the original medical image constructs. Conventional questions ask information about the image from a single dimension and are usually clearly expressed without any confrontational or model-confusing vocabulary, such as those illustrated in Figure 2(b) of the main paper as well as in Figure 10.
@@
Counterfactual Questions: Counterfactual questions present prompts that typically include premises contrary to the factual content of the image. These counterfactual premises often involve incorrect descriptions of various dimensions of a medical image, such as shape, size, location, number, or pathology. The questions are then based on these incorrect descriptions. As illustrated in Figure 2(b) of the main paper and Figure 12, counterfactual questions can be understood as adversarial descriptions designed to test whether the model can detect discrepancies between the prompt and the fundamental facts depicted in the image. This type of question evaluates the model’s ability to search for image features based on textual information and determine whether the model truly understands the image rather than merely fitting the training data domain.
@@ A.5 More Results of Baseline Models in the Med-HallMark
To provide a more comprehensive comparison of different baseline models’ performance on Med-HallMark, we have included additional detailed metrics for various baseline models in both VQA and IRG tasks, as shown in Tables 5 and 7. Additionally, we have provided traditional metrics for each model on counterfactual questions, which are not detailed in the main text, as illustrated in Table 6.
@@ A.6 Limitations
Need for Further Enhancement: Although our hallucination detection model, MediHallDetector, surpasses strong models like GPT-3.5, GPT-4, and Gemini in detection performance and significantly improves efficiency compared to manual evaluation, there remains a gap compared to human preferences. This is due to the inherent limitations of the base models and the intrinsic challenges of LVLMs.
@@ Appendix B Hosting and Maintenance Plan
The authors and corresponding lab will host the dataset and handle maintenance concerns. We have carefully scrutinized the data to ensure that there are no errata for any data points. We have released partial Med-HallMark to GitHub and will be maintaining and updating it continuously.
@@ Appendix C Author Statement
Additionally, all textual content in this work is based on the original medical images. By using these texts, you automatically agree to the license terms of the source datasets. Any violations of the source dataset rules by users will result in them bearing responsibility for any infringement issues.

### v2

@@ 6.3  Evaluations of the MediHallDetector’s Performance
Comparison of different hallucination detection models: We compared our model’s performance with two of the most powerful LLMs: GPT-3.5 and GPT-4 in the context of medical hallucination hierarchy. As depicted in the radar chart in Figure 2(c), the chart illustrates the alignment of detection models with human preferences across six different hallucination levels. GPT-3.5 predominantly classified responses as Critical and Attribute hallucinations, and only correctly identified approximately 1.14% of these instances, indicating a poor understanding of the nuanced definitions of different hallucination levels and an inability to reasonably assess the correctness of $A$ in more complex hallucination classification scenarios. While GPT-4 performed better at detecting the correctness of $A$ in complex prompts, they still struggled with the hierarchical categorization of hallucinations. This was particularly evident in distinguishing prompt-induced hallucinations.
Figure 2(b) compares MedHallDetector with GPT-3.5 and GPT-4 across two dimensions: evaluation time and multi-round evaluation consistency. The results show that GPT models tend to produce varying conclusions in each evaluation, whereas MedHallDetector maintains a high consistency rate of 93.88% across three evaluation rounds. Additionally, MedHallDetector achieves the shortest average inference time per evaluation.
@@ 7  Conclusion
In this paper, we address the challenges of hallucination detection and evaluation in the application of LVLMs in healthcare. We propose a novel benchmark, evaluation metrics, and a detection model specifically designed for the medical domain. In addition to establishing baselines for current mainstream models on the benchmark, we demonstrate the effectiveness of our metrics and model in hallucination evaluation and detection through extensive experimental analysis. We hope this work can significantly improve the reliability of LVLMs in medical applications.
@@ A.1  Details of Med-HallMark
Conventional Medical Questions: In this paper, conventional medical questions refer to the original questions we constructed based on the original medical image constructs. Conventional questions ask information about the image from a single dimension and are usually clearly expressed without any confrontational or model-confusing vocabulary, such as those illustrated in Figure 3 of the main paper as well as in Figure 10.
@@
Counterfactual Questions: Counterfactual questions present prompts that typically include premises contrary to the factual content of the image. These counterfactual premises often involve incorrect descriptions of various dimensions of a medical image, such as shape, size, location, number, or pathology. The questions are then based on these incorrect descriptions. As illustrated in Figure 3 of the main paper and Figure 12, counterfactual questions can be understood as adversarial descriptions designed to test whether the model can detect discrepancies between the prompt and the fundamental facts depicted in the image. This type of question evaluates the model’s ability to search for image features based on textual information and determine whether the model truly understands the image rather than merely fitting the training data domain.
@@ A.5  More Results of Baseline Models in the Med-HallMark
To provide a more comprehensive comparison of different baseline models’ performance on Med-HallMark, we have included additional detailed metrics for various baseline models in both VQA and IRG tasks, as shown in Tables 6 and 8. Additionally, we have provided traditional metrics for each model on counterfactual questions, which are not detailed in the main text, as illustrated in Table 7.
Figure 15: Comparison of different Models on hallucination types.
@@ Appendix B Hosting and Maintenance Plan
The authors and corresponding lab will host the dataset and handle maintenance concerns. We have carefully scrutinized the data to ensure that there are no errata for any data points. We have released the Med-HallMark to GitHub and will be maintaining and updating it continuously.
@@ Appendix D Ethical Statement
All medical images used in this study are sourced from existing open-source datasets. Our use of open-source images strictly adheres to the relevant licensing agreements, and the study involved adding textual annotations to these images to create a new benchmark. Therefore, no new user research was conducted in this study.
For open-source resources, all data are rigorously anonymized and desensitized to avoid ethical issues and do not involve any user studies. Specifically, the protocols for VQA-RAD, MIMIC-CXR, and OpenI are CC0 1.0 Universal, MIT-license, and CC BY-NC-ND 4.0, respectively. Ethical approval was not required as confirmed by the license attached to the open-access data. These protocols allow users to manipulate the data without restriction, including, but not limited to, the right to use, copy, modify, merge, and distribute.
Apart from that, all the studies in this paper were conducted under the supervision of the relevant medical institutions in the host countries and were approved by the ethics committees. Due to the double-blind restriction, we are unable to submit the relevant ethical approvals directly but will seek immediate ethical approval from the ICLR program committee if the article is accepted or if it is required during the review process.